OpenStudy (anonymous):

So very confused........ If the sum of the measures of the interior angles of a convex n-gon is equal to the sum of the measures of the exterior angles, how many sides does the n-gon have?

OpenStudy (dumbcow):

sum of interior angles = 180(n-2) sum of exterior = 360 --> 180(n-2) = 360 solve for n

OpenStudy (anonymous):

Oh, okay I get it now. I just didn't know how to set it up. Thanks!